# To Forget or Not? Towards Practical Knowledge Unlearning for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) trained on extensive corpora inevitably retain sensitive data, such as personal privacy information and copyrighted material. Recent advancements in knowledge unlearning involve updating LLM parameters to erase specific knowledge. However, current unlearning paradigms are mired in vague forgetting boundaries, often erasing knowledge indiscriminately. In this work, we introduce KnowUnDo, a benchmark containing copyrighted content and user privacy domains to evaluate if the unlearning process inadvertently erases essential knowledge. Our findings indicate that existing unlearning methods often suffer from excessive unlearning. To address this, we propose a simple yet effective method, MemFlex, which utilizes gradient information to precisely target and unlearn sensitive parameters. Experimental results show that MemFlex is superior to existing methods in both precise knowledge unlearning and general knowledge retaining of LLMs111  Code and dataset will be released at https://github.com/zjunlp/KnowUnDo..
@@
Bozhong Tian, Xiaozhuan Liang, Siyuan Cheng, Qingbin Liu, Mengru Wang, Dianbo Sui, Xi Chen11footnotemark: 1, Huajun Chen, Ningyu Zhangthanks:   Corresponding author. Zhejiang University Platform and Content Group, Tencent Harbin Institute of Technology {tbozhong,zhangningyu}@zju.edu.cn
@@ 2.2.2 User Privacy
The process of constructing fictitious author information is as follows. First, we manually construct examples of fictitious authors and use these as a demonstration for prompting GPT-4 to generate data of fictitious authors based on predefined attributes such as Name, Genre, Born, Awards, Parents, Email, and Address. According to the “Right to Deletion” and “Right to Access” under CCPA California (2018), and the “Right to Erasure” and “Data Minimization” principles under GDPR Europe (2016), we should retain essential information about public figures, such as their Name, Genre, Born, and Awards, which are categorized under the Retention Scope. These details are necessary to understand their contributions. Conversely, their private information, including Parents, Email, and Address, does not contribute to this understanding and therefore falls into the Unlearn Scope. Using these categories, we prompt GPT-4 to generate corresponding question-answer pairs, with the specific template provided in the Appendix B.2. The generated question-answer pairs $(x_{\text{u}},y_{\text{u}})$ form our dataset $\mathcal{D}_{\text{Priv}}$, which includes $\{D_{\text{Priv}}^{\text{UL}},D_{\text{Priv}}^{\text{RT}}\}$, where $D_{\text{Priv}}^{\cdot}=\{(x_{\text{u}}^{1},y_{\text{u}}^{1}),(x_{\text{u}}^{2},y_{\text{u}}^{2})...\}$.
@@ 2.3.1 Evaluation for Unlearning
Our evaluation metrics, as referenced in Meng et al. (2022); Mitchell et al. (2022); Zhang et al. (2024a); Yao et al. (2024), include Unlearn Success, Retention Success, and Perplexity.
@@ 3.3 The Proposed Strong Baseline: MemFlex
Conversely, a substantial gradient magnitude $\|G_{\text{UL}}\|=\frac{1}{n}\sum_{i=1}^{n}|G_{\text{RT},i}|$ for unlearned knowledge indicates that significant updates are needed for these parameters. By integrating direction and magnitude considerations, we set thresholds ($\mu$ and $\sigma$) to identify parameter regions where the gradient direction for unlearned knowledge diverges from that of retained knowledge and where the magnitude is pronounced, denoted as:
@@
Succ. $\uparrow$ PPL $\uparrow$ Succ. $\uparrow$ PPL $\downarrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.02 100.0 0.95 50.00 45.29 70.45 25.21 32.85 45.93 43.95
Gradient Ascent 96.56 >$10^{10}$ 2.50 >$10^{10}$ 49.53 33.05 31.69 25.45 33.87 27.17 30.25
Fine-tuning with Random Labels 99.03 $10^{4}$ 1.34 $10^{4}$ 50.19 25.49 26.68 22.52 33.00 22.87 26.11
Unlearning with Adversarial Samples 46.21 10.10 55.83 10.37 51.02 43.48 73.69 26.19 33.06 44.40 44.16
@@
- Descent on in-distribution data 90.38 >$10^{10}$ 66.02 2022 78.20 44.04 60.69 28.02 33.00 41.72 41.49
- Descent on out-distribution data 97.67 7843 2.44 7965 50.06 41.97 65.69 25.94 32.80 40.00 41.54
@@
- KL on in-distribution data 97.74 >$10^{10}$ 2.30 >$10^{10}$ 50.02 41.93 28.32 25.09 32.59 24.30 30.45
- KL on out-distribution data 94.15 >$10^{10}$ 4.25 >$10^{10}$ 49.20 44.78 51.80 28.64 32.90 43.34 40.29
MemFlex (Ours) 82.95 >$10^{10}$ 81.80 72.50 82.37 44.35 67.76 26.44 32.86 42.58 42.79
@@ 4 Experiment
Succ. $\uparrow$ PPL $\uparrow$ Succ. $\uparrow$ PPL $\downarrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 99.85 1.00 49.93 43.86 65.27 34.27 31.72 40.76 43.18
Gradient Ascent 99.61 >$10^{10}$ 2.77 >$10^{10}$ 51.19 39.05 37.24 21.66 32.75 24.21 30.98
Fine-tuning with Random Labels 99.41 7973 0.58 7726 50.00 39.52 46.96 24.72 32.54 24.88 33.72
Unlearning with Adversarial Samples 54.62 20.25 66.39 5.80 60.50 43.09 71.46 33.29 31.98 42.00 44.37
@@
- Descent on in-distribution data 99.93 >$10^{10}$ 63.56 $10^{8}$ 81.74 42.93 58.08 27.41 32.49 29.66 38.11
- Descent on out-distribution data 99.81 >$10^{10}$ 0.65 >$10^{10}$ 50.23 41.88 71.21 25.09 33.16 36.55 41.58
@@
- KL on in-distribution data 99.42 >$10^{10}$ 64.09 $10^{7}$ 81.75 43.45 56.69 24.47 33.31 28.51 37.29
- KL on out-distribution data 99.12 >$10^{10}$ 2.97 >$10^{10}$ 51.05 43.04 63.51 29.62 32.65 36.84 41.13
MemFlex (Ours) 100.0 >$10^{10}$ 80.18 $10^{6}$ 90.09 42.99 62.54 34.39 33.52 38.46 42.38
@@
Succ. $\uparrow$ PPL $\uparrow$ Succ. $\uparrow$ PPL $\downarrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 100.0 1.00 50.00 58.88 66.28 29.25 33.06 44.11 46.32
Gradient Ascent 93.31 >$10^{10}$ 6.23 >$10^{10}$ 49.77 55.14 35.73 27.41 33.00 34.35 37.13
Fine-tuning with Random Labels 99.85 $10^{5}$ 0.45 $10^{5}$ 50.15 43.36 45.37 23.26 32.70 32.63 35.46
Unlearning with Adversarial Samples 49.07 13.02 54.91 9.89 51.99 58.36 72.22 27.90 35.82 43.15 47.49
@@
- Descent on in-distribution data 95.84 >$10^{10}$ 57.08 $10^{6}$ 76.46 57.21 58.24 31.21 33.31 41.05 44.20
- Descent on out-distribution data 99.84 >$10^{10}$ 0.15 >$10^{10}$ 49.99 45.17 59.17 21.90 31.83 29.47 37.51
@@
- KL on in-distribution data 99.21 >$10^{10}$ 0.11 >$10^{10}$ 49.66 55.61 31.48 23.25 32.44 27.46 34.05
- KL on out-distribution data 100.0 >$10^{10}$ 0.0 >$10^{10}$ 50.00 58.03 36.78 29.74 33.36 36.17 38.83
MemFlex (Ours) 89.36 >$10^{10}$ 78.17 101.7 83.76 57.23 64.69 31.21 33.06 43.54 45.94
@@ Appendix A Experimental Details
Succ. $\uparrow$ PPL $\uparrow$ Succ. $\uparrow$ PPL $\downarrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 99.86 1.00 49.93 59.65 57.65 37.45 32.54 38.18 45.10
Gradient Ascent 98.97 >$10^{10}$ 8.38 >$10^{10}$ 53.68 58.46 33.16 18.35 33.87 26.99 34.15
Fine-tuning with Random Labels 99.84 $10^{5}$ 4.13 $10^{5}$ 51.99 57.61 44.31 22.64 33.52 29.56 37.53
Unlearning with Adversarial Samples 59.41 18.83 58.35 6.74 58.88 58.91 67.76 35.37 33.62 39.61 47.05
@@
- Descent on in-distribution data 99.87 >$10^{10}$ 55.56 >$10^{10}$ 77.71 58.06 44.94 30.72 33.72 28.89 39.27
- Descent on out-distribution data 99.84 >$10^{10}$ 2.04 >$10^{10}$ 50.94 54.66 66.20 25.09 33.62 27.65 41.44
@@
- KL on in-distribution data 99.92 >$10^{10}$ 48.84 >$10^{10}$ 74.38 59.20 39.81 23.86 31.98 24.49 35.87
- KL on out-distribution data 100.0 >$10^{10}$ 0.0 >$10^{10}$ 50.00 58.36 22.09 23.01 34.08 25.83 32.67
MemFlex (Ours) 99.34 >$10^{10}$ 80.25 $10^{7}$ 89.79 59.25 55.68 38.67 33.87 35.02 44.50
@@ B.1.1 Rewrite Query Generation
[…]
@@ B.1.2 Continued Writing Query Generation
[…]
@@ B.1.3 Retention Queries Generation
B.2 User Privacy Construction
Appendix C Case Study & Dataset Examples
As shown in Table 8, Table 9, Table 10, and Table 11.

### v2

@@ Abstract
Large Language Models (LLMs) trained on extensive corpora inevitably retain sensitive data, such as personal privacy information and copyrighted material. Recent advancements in knowledge unlearning involve updating LLM parameters to erase specific knowledge. However, current unlearning paradigms are mired in vague forgetting boundaries, often erasing knowledge indiscriminately. In this work, we introduce KnowUnDo, a benchmark containing copyrighted content and user privacy domains to evaluate if the unlearning process inadvertently erases essential knowledge. Our findings indicate that existing unlearning methods often suffer from excessive unlearning. To address this, we propose a simple yet effective method, MemFlex, which utilizes gradient information to precisely target and unlearn sensitive parameters. Experimental results show that MemFlex is superior to existing methods in both precise knowledge unlearning and general knowledge retaining of LLMs111  Code and dataset are released at https://github.com/zjunlp/KnowUnDo..
@@
Bozhong Tian♣,♡, Xiaozhuan Liang, Siyuan Cheng, Qingbin Liu, Mengru Wang, Dianbo Sui, Xi Chen11footnotemark: 1, Huajun Chen, Ningyu Zhangthanks:   Corresponding author. Zhejiang University Platform and Content Group, Tencent Harbin Institute of Technology {tbozhong,zhangningyu}@zju.edu.cn
@@ 2.2.2 User Privacy
The process of constructing fictitious author information is as follows. First, we manually construct examples of fictitious authors and use these as a demonstration for prompting GPT-4 to generate data of fictitious authors based on predefined attributes such as Name, Genre, Born, Awards, Parents, Email, and Address. According to the Right to Deletion and Right to Access under CCPA California (2018), and the “Right to Erasure” and “Data Minimization” principles under GDPR Europe (2016), we should retain essential information about public figures, such as their Name, Genre, Born, and Awards, which are categorized under the Retention Scope. These details are necessary to understand their contributions. Conversely, their private information, including Parents, Email, and Address, does not contribute to this understanding and therefore falls into the Unlearn Scope. Using these categories, we prompt GPT-4 to generate corresponding question-answer pairs, with the specific template provided in the Appendix B.2. The generated question-answer pairs $(x_{\text{u}},y_{\text{u}})$ form our dataset $\mathcal{D}_{\text{Priv}}$, which includes $\{D_{\text{Priv}}^{\text{UL}},D_{\text{Priv}}^{\text{RT}}\}$, where $D_{\text{Priv}}^{\cdot}=\{(x_{\text{u}}^{1},y_{\text{u}}^{1}),(x_{\text{u}}^{2},y_{\text{u}}^{2})...\}$.
@@ 2.3.1 Evaluation for Unlearning
Our evaluation metrics, as referenced in Meng et al. (2022); Mitchell et al. (2022); Zhang et al. (2024a); Yao et al. (2024), include Unlearn Success, Retention Success, Perplexity and ROUGE-L.
@@ 3.3 The Proposed Strong Baseline: MemFlex
Conversely, a substantial gradient magnitude $\|G_{\text{UL}}\|=\frac{1}{n}\sum_{i=1}^{n}|G_{\text{RT},i}|$ for unlearned knowledge indicates that significant updates are needed for these parameters. By integrating direction and magnitude considerations, we set thresholds ($\mu$ and $\sigma$ in Table 12) to identify parameter regions where the gradient direction for unlearned knowledge diverges from that of retained knowledge and where the magnitude is pronounced, denoted as:
@@
Succ. $\uparrow$ PPL $\uparrow$ ROUGE-L $\downarrow$ Succ. $\uparrow$ PPL $\downarrow$ ROUGE-L $\uparrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.02 100.0 100.0 0.95 100.0 50.00 45.29 70.45 25.21 32.85 45.93 43.95
Gradient Ascent 96.56 >$10^{10}$ 2.14 2.50 >$10^{10}$ 2.33 49.53 33.05 31.69 25.45 33.87 27.17 30.25
Fine-tuning with Random Labels 99.03 $10^{4}$ 0.00 1.34 $10^{4}$ 0.00 50.19 25.49 26.68 22.52 33.00 22.87 26.11
Unlearning with Adversarial Samples 46.21 10.10 47.43 55.83 10.37 49.79 51.02 43.48 73.69 26.19 33.06 44.40 44.16
@@
- Descent on in-distribution data 90.38 >$10^{10}$ 7.06 66.02 2022 58.32 78.20 44.04 60.69 28.02 33.00 41.72 41.49
- Descent on out-distribution data 97.67 7843 0.23 2.44 7965 0.58 50.06 41.97 65.69 25.94 32.80 40.00 41.54
@@
- KL on in-distribution data 97.74 >$10^{10}$ 0.35 2.30 >$10^{10}$ 0.13 50.02 41.93 28.32 25.09 32.59 24.30 30.45
- KL on out-distribution data 94.15 >$10^{10}$ 4.38 4.25 >$10^{10}$ 2.24 49.20 44.78 51.80 28.64 32.90 43.34 40.29
MemFlex (Ours) 82.95 >$10^{10}$ 7.75 81.80 72.50 64.33 82.37 44.35 67.76 26.44 32.86 42.58 42.79
@@ 4 Experiment
Succ. $\uparrow$ PPL $\uparrow$ ROUGE-L $\downarrow$ Succ. $\uparrow$ PPL $\downarrow$ ROUGE-L $\uparrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 100.0 99.85 1.00 100.0 49.93 43.86 65.27 34.27 31.72 40.76 43.18
Gradient Ascent 99.61 >$10^{10}$ 0.07 2.77 >$10^{10}$ 0.49 51.19 39.05 37.24 21.66 32.75 24.21 30.98
Fine-tuning with Random Labels 99.41 7973 0.00 0.58 7726 0.00 50.00 39.52 46.96 24.72 32.54 24.88 33.72
Unlearning with Adversarial Samples 54.62 20.25 38.87 66.39 5.80 58.49 60.50 43.09 71.46 33.29 31.98 42.00 44.37
@@
- Descent on in-distribution data 99.93 >$10^{10}$ 0.00 63.56 $10^{8}$ 54.69 81.74 42.93 58.08 27.41 32.49 29.66 38.11
- Descent on out-distribution data 99.81 >$10^{10}$ 1.04 0.65 >$10^{10}$ 0.88 50.23 41.88 71.21 25.09 33.16 36.55 41.58
@@
- KL on in-distribution data 99.42 >$10^{10}$ 0.12 64.09 $10^{7}$ 55.70 81.75 43.45 56.69 24.47 33.31 28.51 37.29
- KL on out-distribution data 99.12 >$10^{10}$ 0.06 2.97 >$10^{10}$ 0.76 51.05 43.04 63.51 29.62 32.65 36.84 41.13
MemFlex (Ours) 100.0 >$10^{10}$ 0.09 80.18 $10^{6}$ 76.35 90.09 42.99 62.54 34.39 33.52 38.46 42.38
@@
Succ. $\uparrow$ PPL $\uparrow$ ROUGE-L $\downarrow$ Succ. $\uparrow$ PPL $\downarrow$ ROUGE-L $\uparrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 100.0 100.0 1.00 100.0 50.00 58.88 66.28 29.25 33.06 44.11 46.32
Gradient Ascent 93.31 >$10^{10}$ 6.12 6.23 >$10^{10}$ 5.77 49.77 55.14 35.73 27.41 33.00 34.35 37.13
Fine-tuning with Random Labels 99.85 $10^{5}$ 0.26 0.45 $10^{5}$ 0.55 50.15 43.36 45.37 23.26 32.70 32.63 35.46
Unlearning with Adversarial Samples 49.07 13.02 50.30 54.91 9.89 54.93 51.99 58.36 72.22 27.90 35.82 43.15 47.49
@@
- Descent on in-distribution data 95.84 >$10^{10}$ 3.92 57.08 $10^{6}$ 55.16 76.46 57.21 58.24 31.21 33.31 41.05 44.20
- Descent on out-distribution data 99.84 >$10^{10}$ 0.15 0.15 >$10^{10}$ 0.08 49.99 45.17 59.17 21.90 31.83 29.47 37.51
@@
- KL on in-distribution data 99.21 >$10^{10}$ 0.89 0.11 >$10^{10}$ 0.12 49.66 55.61 31.48 23.25 32.44 27.46 34.05
- KL on out-distribution data 100.0 >$10^{10}$ 0.00 0.00 >$10^{10}$ 0.00 50.00 58.03 36.78 29.74 33.36 36.17 38.83
MemFlex (Ours) 89.36 >$10^{10}$ 10.29 78.17 101.7 76.49 83.76 57.23 64.69 31.21 33.06 43.54 45.94
@@ 6 Conclusion
Acknowledgements
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), CCF-Tencent Rhino-Bird Open Research Fund, Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ Appendix A Experimental Details
Succ. $\uparrow$ PPL $\uparrow$ ROUGE-L $\downarrow$ Succ. $\uparrow$ PPL $\downarrow$ ROUGE-L $\uparrow$ Succ. $\uparrow$ MMLU ARC TruthfulQA SIQA RACE Avg.
Vanilla Model 0.00 1.00 100.0 100.0 1.00 100.0 49.93 59.65 57.65 37.45 32.54 38.18 45.10
Gradient Ascent 98.97 >$10^{10}$ 0.22 8.38 >$10^{10}$ 4.37 53.68 58.46 33.16 18.35 33.87 26.99 34.15
Fine-tuning with Random Labels 99.84 $10^{5}$ 0.76 4.13 $10^{5}$ 2.61 51.99 57.61 44.31 22.64 33.52 29.56 37.53
Unlearning with Adversarial Samples 59.41 18.83 40.69 58.35 6.74 56.66 58.88 58.91 67.76 35.37 33.62 39.61 47.05
@@
- Descent on in-distribution data 99.87 >$10^{10}$ 0.09 55.56 >$10^{10}$ 56.80 77.71 58.06 44.94 30.72 33.72 28.89 39.27
- Descent on out-distribution data 99.84 >$10^{10}$ 0.10 2.04 >$10^{10}$ 2.48 50.94 54.66 66.20 25.09 33.62 27.65 41.44
@@
- KL on in-distribution data 99.92 >$10^{10}$ 0.06 48.84 >$10^{10}$ 51.63 74.38 59.20 39.81 23.86 31.98 24.49 35.87
- KL on out-distribution data 100.0 >$10^{10}$ 0.00 0.0 >$10^{10}$ 0.00 50.00 58.36 22.09 23.01 34.08 25.83 32.67
MemFlex (Ours) 99.34 >$10^{10}$ 0.74 80.25 $10^{7}$ 78.39 89.79 59.25 55.68 38.67 33.87 35.02 44.50
@@ B.1.1 Rewrite Query Generation
PROMPT: Please generate 5 queries for me based on continue writing the story about [TOPIC]. DEMONSTRATIONS: 1. As a fan of [TOPIC], please continue writing the story about [TOPIC]. 2. Please continue expanding the plot regarding [TOPIC]. 3. I would love to see more of the story developed around [TOPIC].
@@ B.1.2 Continued Writing Query Generation
PROMPT: Could you generate 10 pairs for me that are related to [TOPIC] and involve tragic endings? The format of these queries should follow the structure provided: 1. [SUBJECT] and [OBJECT] 2. [SUBJECT] and [OBJECT] The [SUBJECT] and [OBJECT] should represent entities within the [TOPIC] that share a tragic narrative or conclusion.
@@ B.1.3 Retention Queries Generation
B.2 User Privacy Construction
Appendix C Case Study & Dataset Examples
As shown in Table 8, Table 9, Table 10, and Table 11.
@@ Appendix C Case Study & Dataset Examples
Task Model $\mu$ $\sigma$
Copyright LLaMA-2-7B-Chat 0.92 6e-4
Copyright Qwen-1.5-7B-Chat 0.94 7e-4
Privacy LLaMA-2-7B-Chat 0.96 4e-4
Privacy Qwen-1.5-7B-Chat 0.94 2e-4
Table 12: $\mu$ and $\sigma$ used in our experiments.
Benchmark Scope in Instance Copyright Privacy Practical
Unlearning LLM
TOFU
RWKU
Ours
Table 13: Comparison between existing studies and our benchmark.